40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11134     Entry time: Wed Mar 11 19:15:03 2015
Author: Koji 
Type: Summary 
Category: LSC 
Subject: ROUGH calibration of the darm spectrum during the full PRFPMI lock 

I made very rough calibration of the DARM spectra before and after the transition for the second lock on Mar 8.

The cavity pole (expected to be 4.3kHz) was not compensated. Also the servo bump was not compensated.

[Error calibration]

While the DARM/CARM were controlled with ALS, the calibration of them are provided by the ALS phase tracker calibration.
i.e 1 degree = 19.23kHz

This means that the calibration factor is

DARM [deg] * 19.23e3 [Hz/deg] / c [m Hz] * lambda [m] * L_arm [m]
= DARM* 19.23e3/299792458*1064e-9*38.5 = 2.6e-9 *DARM [m]

[Feedback calibration]

Then, the feedback signal was calibrated by the suspension response (f=1Hz, Q=5)
so that the error and feedback signals can match at 100Hz.

This gave me the DC factor of 5e-8.


The spectra at 1109832200 (ALS only, even not on the resonance) and 1109832500 (after DARM/CARM transitions) were taken.
Jenne said that the whitening filters for AS55Q was not on.

Attachment 1: 150308_DARM_ROUGH_CALIB.pdf  120 kB  Uploaded Wed Mar 11 20:25:26 2015  | Hide | Hide all
150308_DARM_ROUGH_CALIB.pdf
ELOG V3.1.3-